In a package of candy coated chocolates, 30 percent of the candies are brown, 20% of them are yellow, and 20% of them are red, there are 10% each of orange, green,and blue. if Ronan reaches into a package and pulls out 12 candies, what is the probability that
A) Exactly 2 of them would be red
B) at most 2 of them would be red
C) none of them would be blue
Round your answers to the nearest tenth of a percent.

Answer by stanbon(57290)

In a package of candy coated chocolates, 30 percent of the candies are brown, 20% of them are yellow, and 20% of them are red, there are 10% each of orange, green,and blue. if Ronan reaches into a package and pulls out 12 candies, what is the probability that
A) Exactly 2 of them would be red
P(2red of 12) = 12C2(0.20)^2(0.80)^10 = 0.2835
------------
B) at most 2 of them would be red
Using a Ti calculator you get:
P(0,1,or 2 are red) = binomcdf(12,0.2,2) = 0.5583
---------------
C) none of them would be blue
P(12 not blue) = 0.9^12 = 0.2824
---------------------
Cheers,
Stan H.